OpenStudy (anonymous):

2 cos theta = 1, find all solutions

OpenStudy (anonymous):

Step by step help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so think of theta as a variable

OpenStudy (anonymous):

if theta was an 'x' how would you solve?

OpenStudy (anonymous):

to be honest I have no clue.

OpenStudy (shaik0124):

2 cos theta=1 cos theta=1/2 theta=cos inverse(1/2) theta=60degrees

OpenStudy (anonymous):

ok, so i would divide 2 from both sides?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):